Three persons Alice, Belen and Carol, played games. At the beginning of the game,the money they had were in the ratio 3:2:1. At the end of the game, their money were in the ratio 11:8:5. Who won and who lost?

Answer by ankor@dixie-net.com(22740)

Three persons Alice, Belen and Carol, played games
. At the beginning of the game,the money they had were in the ratio 3:2:1.
At the end of the game, their money were in the ratio 11:8:5.
Who won and who lost?
;
One way is make an assumption of actual amts for each, that reflect the given ratio
Beginning
A has $30
B has $20
C has $10
----------
Total:$60
:
After the game the total will still be $60
let x = the multiplier
A has 11x
B has 8x
C has 5x
:
Find the multiplier x
11x + 8x + 5x = 60
24x = 60
x = 60/24
x = 2.5 is the multiplier
then
11*2.5 = $27.50, A lost $2.50
8* 2.5 = $20.00, B came out even
5* 2.5 = $12.50, C gained $2.50
--------------------
Check: $60.00, total remains the same